# Analytical Insights into Constant-Roll Condition: Extending the Paradigm to Non-Canonical Models

### v1

@@ 4.1 Field-Constant-Roll Condition
We now discuss the dynamics of the constant propagation class (4.1) constrained with the FCR condition (3.7). In this case, if $\mathcal{F}(\phi)$ is known, one can start the study by integrating the constant-roll condition (3.7) written in the following form:
@@
where equation (4.3) is implemented. Having $\mathcal{F}(\phi)$, one can solve this equation to find $H(\phi)$. Once Hubble parameter has been determined, the field velocity can be derived using equation (4.3) as follows:
@@
and utilizing equation (4.2) for the potential we easily find
@@
Note that, contrary to the potential expression (4.10), these relations are not directly dependent on $\mathcal{F}$ function. It is clear from equation (4.11) that the derivative of effective potential vanishes at zeros of $H$ and $\dot{\phi}$. On the other hand zeros of $H_{,\phi}$ and $\dot{\phi}$ coincide (see (4.9)) and we find the effective potential extremas at zeros of $H$ and ${H}_{,\phi}$. Based on equation (4.12), we notice that, like the canonical case, at extrema the concavity of effective potential is determined by $\beta_{\text{\tiny{F}}}$ values; the effective potential is concave for $\beta_{\text{\tiny{F}}}<-3c_{s}^{2}$ and $\beta_{\text{\tiny{F}}}>0$, and it is convex for $-3c_{s}^{2}<\beta_{\text{\tiny{F}}}<0$. For $\beta_{\text{\tiny{F}}}=-3c_{s}^{2}$ the effective potential is flat.
@@
$N-N_{0}=\frac{c_{s}^{2}}{\beta_{\text{\tiny{F}}}}\ln\left(\frac{H_{,\phi}}{\mathcal{F}}\right),$ (4.13)
We assume that $H_{,\phi}$ has a zero at $\phi=\tilde{\phi}$. From equation (4.13) we notice that as $\phi\rightarrow\tilde{\phi}$, the expression $\ln\left(H_{,\phi}/\mathcal{F}\right)$ diverges and $\beta_{\text{\tiny{F}}}\left(N-N_{0}\right)$ goes to negative infinity, if $\mathcal{F}(\tilde{\phi})\neq 0$. Using equations (2.4), (4.3), and (4.13) we can find the following general relations for the first and second slow-roll parameters:
@@
As the scalar field starts to traverse the extrema of the effective potential at $\phi=\tilde{\phi}$, $\mathcal{F}(\phi)$ and $H(\phi)$ approach the constant values of $\mathcal{F}(\tilde{\phi})$ and $H(\tilde{\phi})$ and we have an exponentially decreasing (increasing) $\epsilon_{1}$ for positive (negative) $\mathcal{F}(\tilde{\phi})$. Furthermore, if the time scale of $\mathcal{F}$ variation is large enough and $\epsilon_{1}$ is sufficiently small, for the second slow-roll parameter we get $\epsilon_{2}\propto 2\lambda\beta_{\text{\tiny{F}}}$. Therefore, the model exhibits a quasi-canonical behavior as $\phi\rightarrow\tilde{\phi}$ (and if we have $\mathcal{F}(\tilde{\phi})=1$, then we will observe normal canonical behavior). if this phase is characterized by a large negative $\tilde{\beta}_{\text{\tiny{F}}}\equiv\lambda\beta_{\text{\tiny{F}}}<-3/2$, its dynamics enter the non-attractor era lasting some $e$-folds of evolution during which the first slow-roll parameter decays exponentially.
The solutions with $\beta_{\text{\tiny{F}}}<-3c_{s}^{2}$ and $\beta>0$ are of phenomenological interest. In this case the scalar field rolls up a bump-like region and asymptotically comes to rest at the top of the bump. This will be the case for the fine-tuned initial condition and these constant-roll solutions lead classically to eternal inflation. Under quantum perturbations, however, this will be a transient solution As we will show in an example in section 4.3, the generic attractor solution is the dual case $\beta_{\text{\tiny{F}}}\rightarrow-(\beta_{\text{\tiny{F}}}+3c_{s}^{2})$ The transition from the smaller $\beta_{\text{\tiny{F}}}$ to the larger value $-(\beta_{\text{\tiny{F}}}+3c_{s}^{2})$ can be seen through the Klein-Gordon equation (4.4) linearized around the maximum of the effective potential,
$\ddot{\phi}+3c_{s}^{2}H(\tilde{\phi})\dot{\phi}-\beta_{\text{\tiny{F}}}(\beta_{\text{\tiny{F}}}+3c_{s}^{2})H^{2}(\tilde{\phi})(\phi-\tilde{\phi})\approx 0.$ (4.15)
which is invariant under the duality transform. By utilizing (4.11), an equivalent definition of $\beta_{\text{\tiny{F}}}$ is given by $\beta_{\text{\tiny{F}}}=-3c_{s}^{2}+\mathcalboondox{V}_{,\phi}/(H\dot{\phi})$. This form indicates that (if the field velocity is large enough so that the field can pass the extrema), when the field evolves through the critical point $\phi=\tilde{\phi}$, the parameter $\beta_{\text{\tiny{F}}}$ is forced to $-3c_{s}^{2}$. The value of $-3c_{s}^{2}$ for the $\beta_{\text{\tiny{F}}}$ at the maximum of the effective potential is a feature of FCR condition, and as we will show later, it will depend on the $\mathcal{F}$ profile in the HCR (and equivalently PCR) models.
@@ 4.2 Hubble-Constant-Roll and Potential-Constant-Roll Conditions
We discussed that in the constant propagation class (4.1) the HCR and PCR conditions are analogous. Therefore, we only consider HCR conditions (3.9) here. We know that the Hubble parameter in these models are similar to the canonical case. Accordingly, the first and second slow-roll parameters are given by
@@
We observe that in general, slow-roll parameters here differ from (4.14) parameters derived using the FCR condition. However, at the critical point $\tilde{\phi}$, both HCR and FCR constrained models exhibit an approximately identical quasi-canonical behavior with $\beta_{\text{\tiny{H}}}=\tilde{\beta}_{\text{\tiny{F}}}=\lambda\beta_{\text{\tiny{F}}}$. Indeed, this is well consistent with our previous general argument that in constant propagation models three constant-roll conditions are homologous when the timescale of $\mathcal{F}$ variation is large (see equation (4.7)).
After removing $\mathcal{F}$ between equations (4.2) and (4.3), we will have
@@
In this equation, we can write the time derivative in terms of $N$ and use Hubble parameters (2.13), (2.14), (2.15), and (2.29) to find the following potentials respectively:
@@
and integrating it out to find $H(\phi)$. The field is then calculated using the relation $H(\phi)=H(t)=H(N)$.
The effective potential gradient can be obtained from the general expression (3.6) and HCR condition (3.9) in the form of
@@
$\mathcalboondox{V}_{,\phi\phi}\approx-c_{s}^{4}H^{2}\beta_{\text{\tiny{V}}}\left(3+\beta_{\text{\tiny{V}}}\right).$ (4.24)
It is clear that at the critical point $\tilde{\phi}$, where $\mathcal{F}$ is roughly constant, expressions (4.23) and (4.24) become identical to those in the FCR constrained models, (4.11) and (4.12), with $\beta_{\text{\tiny{H}}}=\lambda c_{s}^{2}\beta_{\text{\tiny{V}}}=\lambda\beta_{\text{\tiny{F}}}$. Therefore, based on our analysis, we expect to encounter the duality between $\beta_{\text{\tiny{H}}}$ and $-(\beta_{\text{\tiny{H}}}+3\lambda c_{s}^{2})$ solutions. Nevertheless, this discussion does not show that the model will actually transition to the dual solution. To determine whether $\beta_{\text{\tiny{H}}}$ is parametrically forced to $-(\beta_{\text{\tiny{H}}}+3\lambda c_{s}^{2})$, we need to conduct a stability analysis that considers perturbations in the initial values of the model. We will perform such an analysis in the example provided in section 4.3. Indeed, approximation (4.15) only holds at the beginning and end of the transition trajectory, where the model approximately follows a constant-roll scenario. Hence, the transition trajectory cannot be accurately described by this approximation. The form of the transition trajectory generally depends on the specific form of the action. As a simple demonstration of this dependence, we can use equation (4.23) to express the constant-roll parameter as $\beta_{\text{\tiny{V}}}=\dot{\mathcal{F}}/(2\lambda c_{s}^{2}H\mathcal{F})-\mathcalboondox{V}_{,\phi}/(c_{s}^{2}H\dot{\phi})+3$. This relation highlights that the field velocity at the extrema of the effective potential, where $\mathcalboondox{V}_{,\phi}=0$, depends on the specific form of the parameter $\mathcal{F}$ (unlike in the FCR case, where it is always equal to $-3c_{s}^{2}$ irrespective of the form of the $\mathcal{F}$ function).
Using general relations (3.9) and (4.3), number of $e$-folds and the field are related by
@@ 4.3 An Example: k/G Inflation
where we have chosen $c_{s}=\lambda=1$. These models are well motivated and studied in [67]. For this reasoning, we found it interesting to study its application in constant-roll inflationary models. For $c_{s}=1$ general relations (4.8) and (4.22) become
@@
ptRefer to caption    Refer to captionptRefer to caption     Refer to caption
@@
Figure 4: The plots of the Hubble parameter (4.31) (top left panel), number of $e$-folds (4.32) (top right panel), potential (4.33) and effective potential (4.35) (bottom left panel), and slow-roll parameters (4.36) and (4.37) (bottom right panel), once for $\beta_{\text{\tiny{F}}}=-0.5$ and $\gamma=1$ (solid lines) and once for $\beta_{\text{\tiny{F}}}=0.5$ and $\gamma=-1$ (dashed lines). The parameter $\mathcalboondox{V}_{0}$ in effective potential is taken to be $1.475$. Two vertical grid lines in all panels correspond to two $H_{,\phi}$ zeros at $\phi_{\text{\tiny{F}}}=0$ and $\phi_{\text{\tiny{F}}}=a_{1}\approx-2.34$.
Let us first examine the FCR solution (4.29). For convenience we consider a particular case with $C_{2}=0$ as
$H(\phi_{\text{\tiny{F}}})=MAi^{\prime}(\phi_{\text{\tiny{F}}}),$ (4.31)
where we have defined $M\equiv C_{1}$, $\bar{\beta}_{\text{\tiny{F}}}\equiv\sqrt[3]{-\beta_{\text{\tiny{F}}}\gamma/2}$, and $\phi_{\text{\tiny{F}}}\equiv\bar{\beta}_{\text{\tiny{F}}}\phi$. A plot of the Hubble parameter (4.31) can be seen in Figure 4 (top left panel). To have a real Hubble parameter, we need to choose $\gamma\beta_{\text{\tiny{F}}}<0$. As we wish to examine the model for both positive and negative $\beta_{\text{\tiny{F}}}$, we choose $\beta_{\text{\tiny{F}}}=-0.5$ with $\gamma=1$, as well as $\beta_{\text{\tiny{F}}}=0.5$ with $\gamma=-1$. These two cases share the same Hubble parameter but have different dynamics. Note that in our choices $\bar{\beta}_{\text{\tiny{F}}}$ is positive and zeros of $H_{,\phi}\propto\phi_{\text{\tiny{F}}}Ai(\phi_{\text{\tiny{F}}})$ are located at $\phi_{\text{\tiny{F}}}=0$ and isolated values of $\phi_{\text{\tiny{F}}}$ on negative real axis. Following the notation of reference [77], we show $s^{\text{th}}$ zeros of $Ai(z)$ and $Ai^{\prime}(z)$ with $a_{s}$ and $a^{\prime}_{s}$, respectively [^3] Here we only analyse the model around $a_{1}$, $a^{\prime}_{1}\approx-1.02$, and $a^{\prime}_{2}\approx-3.25$ points.. In particular, we will discuss the model around $\phi_{\text{\tiny{F}}}=0$ and $\phi_{\text{\tiny{F}}}=a_{1}\approx-2.34$ points where $H_{,\phi}$ becomes zero and we expect quasi-canonical behavior. It is important to note that at $\phi_{\text{\tiny{F}}}=0$ the Hubble parameter is negative, which means the universe is collapsing and is not acceptable. However, we can easily avoid this problem by choosing a negative $M$.
@@
The plot of $N(\phi_{\text{\tiny{F}}})$ can be seen in the top right panel of Figure 4. We see that $N(\phi_{\text{\tiny{F}}})$ is singular at $\phi_{\text{\tiny{F}}}=a_{1}$, where $H_{,\phi}\rightarrow 0$, whereas at $\phi_{\text{\tiny{F}}}=0$ it is continuous and finite. It is because $\mathcal{F}=\gamma\phi$ also tends to zero at $\phi_{\text{\tiny{F}}}=0$, making $H_{,\phi}/\mathcal{F}$ non-zero (see equation (4.13)). As a result, we expect quasi-canonical behavior at $\phi_{\text{\tiny{F}}}=a_{1}$ but not at $\phi_{\text{\tiny{F}}}=0$.
@@
ptRefer to caption     Refer to caption
@@
Figure 5: Another plot of the number of $e$-folds (4.32) (left panel), potential (4.33) (solid lines in right panel), and effective potential (4.35) (dashed lines in right panel) for $\gamma=1$ and $\mathcalboondox{V}_{0}=1.475$. Vertical grid line at $\phi_{\text{\tiny{F}}}=a_{1}$ corresponds to singularities of $N(\phi_{\text{\tiny{F}}})$ functions.
where $\mathcalboondox{V}_{0}$ is the constant of integration. Potential (4.33) and effective potential (4.35) are plotted in Figure 4 (bottom left panel). The effective potential shows a roller-coaster profile with plateau-like regions connected to steep cliffs. Effective potential extrema are located at $a_{s}$ and $a^{\prime}_{s}$ which correspond to zeros of $H_{,\phi}$ and $H$, respectively. We recall that in FCR models the effective potential is not directly proportional to $\mathcal{F}$ and its concavity at extrema depends only on the sign of the parameter $\beta_{\text{\tiny{F}}}$. This is not the case for the potential. In Figure 4 we see that for $\beta_{\text{\tiny{F}}}=-0.5$ the effective potential is convex around $\phi_{\text{\tiny{F}}}=a_{1}$ but for $\beta_{\text{\tiny{F}}}=0.5$ it is concave. However, potential is concave in both cases.
Depending on the field initial value, one can think of different inflationary scenarios. In the case of $\beta_{\text{\tiny{F}}}=-0.5$, the singularity of $N(\phi)$ plot at $a_{1}$ (and similarly at $a_{s}$) shows that if the analytic field (that is described by constant $\beta_{\text{\tiny{F}}}$) starts at the neighbourhood of $\phi_{\text{\tiny{F}}}=a_{1}$, it will eventually rest at $a_{1}$. For $\phi_{\text{\tiny{F}}}>a^{\prime}_{1}$, however, the field moves toward the positive infinity. The field in $\beta_{\text{\tiny{F}}}=0.5$ case always rolls in the opposite direction to the $\beta_{\text{\tiny{F}}}=-0.5$ case. It is important to note that $\mathcal{F}=0$, which in our example happens at $\phi_{\text{\tiny{F}}}=0$, is the singular point of the FCR condition (4.27), which means our solutions are not viable around this point and the field cannot surpass it.
Next we derive and analyze slow-roll parameters. Using definition (2.4) and equations (4.9) and (4.31), for the first and second slow-roll parameters we find
@@
Figure 4 (bottom right panel) illustrates these parameters. At $\phi_{\text{\tiny{F}}}=a_{1}$, $\epsilon_{1}$ is approaching zero, while $\epsilon_{2}$ is roughly $2\beta_{\text{\tiny{F}}}$. When $\beta_{\text{\tiny{F}}}=-0.5$, the field moves toward the extrema at $a_{1}$, where we observe quasi-canonical behavior, whereas when $\beta_{\text{\tiny{F}}}=0.5$, the field moves away from $a_{1}$ point, which means we gradually leave the quasi-canonical phase. When $\beta_{\text{\tiny{F}}}$ is positive, $\epsilon_{1}$ grows in time and we can find points where the first slow-roll parameter reaches 1 and inflation ends. Additionally, since we have $\epsilon_{2}\approx 2\beta_{\text{\tiny{F}}}$ when $\phi_{\text{\tiny{F}}}\rightarrow a_{1}$, choosing $\beta_{\text{\tiny{F}}}<-1$ can potentially lead to a burst in the power spectrum and the formation of primordial black holes. We will talk more about this mechanism in another example in section 5.3.
@@
In Figure 4 we showed how the concavity of potential and effective potential changes around $\beta_{\text{\tiny{F}}}=0$. In Figure 5 (right panel) we plotted the potential and effective potential again, this time for $\beta_{\text{\tiny{F}}}=-2$, $-3$, and $-4$, around the critical point $\phi_{\text{\tiny{F}}}=a_{1}$, to show how the concavity changes around $\beta_{\text{\tiny{F}}}=-3$. We can see that the potential is convex for $\beta_{\text{\tiny{F}}}=-4$ and concave for $\beta_{\text{\tiny{F}}}=-2$. In each case, since $\mathcal{F}$ is negative, the concavity of effective potential is in the opposite direction of the potential. When $\beta_{\text{\tiny{F}}}=-3$, we have $\mathcalboondox{V}_{,\phi\phi}=\mathcalboondox{V}_{,\phi}=0$ and effective potential is flat throughout. There is, however, only one point (at the singularity of $N(\phi_{\text{\tiny{F}}})$) where the first and second derivatives of potential are zero. Actually, we do not have a flat USR like potential for any choice of the constant-roll parameter $\beta_{\text{\tiny{F}}}$. To study the USR limit of this model, one should use PCR condition or equivalently HCR condition.

### v3

@@ 4.1 Field-Constant-Roll Condition
We now discuss the dynamics of the constant propagation class of models (4.1) constrained with the FCR condition (3.7). In this case, if $\mathcal{F}(\phi)$ is known, one can start the study by integrating the constant-roll condition (3.7) written in the following form:
@@
where equation (4.3) is being put to use in the above relation. Having $\mathcal{F}(\phi)$, one can solve this equation to find $H(\phi)$. Once the Hubble parameter has been determined, the field velocity can be derived using equation (4.3) as follows:
@@
and by utilizing equation (4.2) for the potential we easily find
@@
Note that, contrary to the potential expression (4.10), these relations are not directly dependent on $\mathcal{F}$ function. It is clear from equation (4.11) that the derivative of effective potential vanishes at zeros of $H$ and $\dot{\phi}$. On the other hand zeros of $H_{,\phi}$ and $\dot{\phi}$ coincide (see (4.9)) and we find the effective potential extrema at zeros of $H$ and ${H}_{,\phi}$. Based on equation (4.12), we notice that, like the canonical case, at extrema the concavity of effective potential is determined by $\beta_{\text{\tiny{F}}}$ values; it is concave for $\beta_{\text{\tiny{F}}}<-3c_{s}^{2}$ and $\beta_{\text{\tiny{F}}}>0$, and is convex for $-3c_{s}^{2}<\beta_{\text{\tiny{F}}}<0$. For $\beta_{\text{\tiny{F}}}=-3c_{s}^{2}$ the effective potential is flat.
@@
$N-N_{0}=\frac{c_{s}^{2}}{\beta_{\text{\tiny{F}}}}\ln\left(\frac{H_{,\phi}}{\mathcal{F}}\right).$ (4.13)
Let us assume that $H_{,\phi}$ has a zero at $\phi=\tilde{\phi}$. From equation (4.13) we notice that as $\phi\rightarrow\tilde{\phi}$, the expression $\ln\left(H_{,\phi}/\mathcal{F}\right)$ diverges and $\beta_{\text{\tiny{F}}}\left(N-N_{0}\right)$ goes to negative infinity, if $\mathcal{F}(\tilde{\phi})\neq 0$. Using equations (4.3) and (4.13) we can find the following general relations for the first and second slow-roll parameters:
@@
As the scalar field moves toward the extrema of the effective potential at $\phi=\tilde{\phi}$, $\mathcal{F}(\phi)$ and $H(\phi)$ approach the constant values of $\mathcal{F}(\tilde{\phi})$ and $H(\tilde{\phi})$ and we have an exponentially decreasing (increasing) $|\epsilon_{1}|$ for positive (negative) $\beta_{\text{\tiny{F}}}$. Furthermore, if the time scale of $\mathcal{F}$ variation is large enough (so that we can neglect the last term in $\epsilon_{2}$ expression) and $\epsilon_{1}$ is sufficiently small, for the second slow-roll parameter we get $\epsilon_{2}\approx 2\lambda\beta_{\text{\tiny{F}}}$. Therefore, the model exhibits a quasi-canonical behavior in the vicinity of $\tilde{\phi}$ (in the sense that at this limit slow-roll parameters behave like the canonical case while the action is not canonical. If we have $\mathcal{F}(\tilde{\phi})=1$, then we will observe normal canonical behavior). If this phase is characterized by a large negative $\tilde{\beta}_{\text{\tiny{F}}}\equiv\lambda\beta_{\text{\tiny{F}}}<-3/2$, its dynamics enter the non-attractor era lasting some $e$-folds of evolution during which the first slow-roll parameter decays exponentially.
The solutions with $\beta_{\text{\tiny{F}}}<-3c_{s}^{2}$ are of phenomenological interest. In this case, the scalar field rolls up a bump-like region and asymptotically comes to rest at the top of the bump. This will be the case for the fine-tuned initial condition and these constant-roll solutions lead classically to eternal inflation. Under quantum perturbations, however, this will be a transient solution. As we will show in an example in section 4.3, the generic attractor solution is the dual case $\beta_{\text{\tiny{F}}}\rightarrow-(\beta_{\text{\tiny{F}}}+3c_{s}^{2})$ (which for $c_{s}=1$, it gives rise to the duality relation discussed in [17, 21] for canonical case). The transition from the smaller $\beta_{\text{\tiny{F}}}$ to the larger value $-(\beta_{\text{\tiny{F}}}+3c_{s}^{2})$ can be seen through the Klein-Gordon equation (4.4) linearized around the maximum of the effective potential,
$\ddot{\phi}+3c_{s}^{2}H(\tilde{\phi})\dot{\phi}-\beta_{\text{\tiny{F}}}(\beta_{\text{\tiny{F}}}+3c_{s}^{2})H^{2}(\tilde{\phi})(\phi-\tilde{\phi})\approx 0,$ (4.15)
which is invariant under the duality transform. By utilizing (4.11), an equivalent definition of $\beta_{\text{\tiny{F}}}$ is given by $\beta_{\text{\tiny{F}}}=-3c_{s}^{2}-\mathcalboondox{V}_{,\phi}/(H\dot{\phi})$. This form indicates that (if the field velocity is large enough so that the field can pass the extrema) when the field evolves through the critical point $\phi=\tilde{\phi}$, the parameter $\beta_{\text{\tiny{F}}}$ is forced to $-3c_{s}^{2}$. The value of $-3c_{s}^{2}$ for the $\beta_{\text{\tiny{F}}}$ at the maximum of the effective potential is a feature of FCR condition, and as we will show later, it will depend on the $\mathcal{F}$ profile in the HCR (and equivalently PCR) models.
@@ 4.2 Hubble-Constant-Roll and Potential-Constant-Roll Conditions
We discussed that in the constant propagation class (4.1) the HCR and PCR conditions are analogous. Therefore, we only consider HCR condition (3.9) here. The Hubble parameter in these models is similar to the canonical case. Accordingly, the first and second slow-roll parameters are given by
@@
We observe that in general slow-roll parameters here differ from (4.14) parameters derived using the FCR condition. However, at the critical point $\tilde{\phi}$, both HCR and FCR constrained models exhibit an approximately identical quasi-canonical behavior with $\beta_{\text{\tiny{H}}}=\tilde{\beta}_{\text{\tiny{F}}}\equiv\lambda\beta_{\text{\tiny{F}}}$. Indeed, this is well consistent with our previous general argument that in constant propagation models three constant-roll conditions are homologous when the timescale of $\mathcal{F}$ variation is large (see equation (4.7)).
After removing $\mathcal{F}$ between equations (4.2) and (4.3) for the potential will have
@@
In this equation, we can write the time derivative in terms of $N$ and use Hubble parameters (2.13)-(2.15), and (2.29) to find the following potentials respectively:
@@
and integrate it out to find $H(\phi)$. The field is then calculated using the relation $H(\phi)=H(t)=H(N)$.
The effective potential gradient can be obtained from the general expression (3.6) and the HCR condition (3.9) in the form of
@@
$\mathcalboondox{V}_{,\phi\phi}\approx-c_{s}^{4}H^{2}(\tilde{\phi})\beta_{\text{\tiny{V}}}\left(3+\beta_{\text{\tiny{V}}}\right).$ (4.24)
This term appears in the linearized Klein-Gordon equation around $\phi=\tilde{\phi}$. It is clear that at the critical point $\tilde{\phi}$ expressions (4.23) and (4.24) become similar to those in the FCR constrained models, (4.11) and (4.12), with $\beta_{\text{\tiny{H}}}=\lambda c_{s}^{2}\beta_{\text{\tiny{V}}}=\lambda\beta_{\text{\tiny{F}}}$. Therefore, based on our analysis, we expect to encounter the duality between $\beta_{\text{\tiny{H}}}$ and $-(\beta_{\text{\tiny{H}}}+3\lambda c_{s}^{2})$ solutions. Nevertheless, this discussion does not show that the model will actually transition to the dual solution. To determine whether $\beta_{\text{\tiny{H}}}$ is parametrically forced to $-(\beta_{\text{\tiny{H}}}+3\lambda c_{s}^{2})$, we need to conduct a stability analysis that considers perturbations in the initial values of the model. We will perform such an analysis in the example provided in section 4.3. Indeed, approximation (4.15) only holds at the beginning and end of the transition trajectory, where the model approximately follows a constant-roll scenario. Hence, the transition trajectory cannot be accurately described by this approximation. The form of the transition trajectory generally depends on the specific form of the action. As a simple demonstration of this dependence, we can use equation (4.23) to express the constant-roll parameter as $\beta_{\text{\tiny{V}}}=\dot{\mathcal{F}}/(2\lambda c_{s}^{2}H\mathcal{F})-\mathcalboondox{V}_{,\phi}/(c_{s}^{2}H\dot{\phi})+3$. This relation highlights that at the extrema of the effective potential, where $\mathcalboondox{V}_{,\phi}=0$, parameter $\beta_{\text{\tiny{V}}}$ depends on the specific form of the parameter $\mathcal{F}$ (unlike in the FCR case, where it is always equal to $-3c_{s}^{2}$ irrespective of the form of the $\mathcal{F}$ function and field velocity).
Using relations (3.9) and (4.3), number of $e$-folds and the field are related by
@@ 4.3 An Example: k/G Inflation
where we have chosen $c_{s}=\lambda=1$. These models are well motivated and studied in [68]. For this reason, we found it interesting to study its application in constant-roll inflationary models. For $c_{s}=1$ general relations (4.8) and (4.22) become
@@
Let us first examine the FCR solution (4.29). For convenience, we consider a particular case with $C_{2}=0$ as
$H(\phi_{\text{\tiny{F}}})=MAi^{\prime}(\phi_{\text{\tiny{F}}}).$ (4.31)
Here, we have defined $M\equiv C_{1}$, $\bar{\beta}_{\text{\tiny{F}}}\equiv\sqrt[3]{-\beta_{\text{\tiny{F}}}\gamma/2}$, and $\phi_{\text{\tiny{F}}}\equiv\bar{\beta}_{\text{\tiny{F}}}\phi$. A plot of the Hubble parameter (4.31) is presented in Figure 4 (top left panel). To have a real Hubble parameter, we need to choose $\gamma\beta_{\text{\tiny{F}}}<0$. As we wish to examine the model for both positive and negative $\beta_{\text{\tiny{F}}}$, we choose $\beta_{\text{\tiny{F}}}=-0.5$ with $\gamma=1$, as well as $\beta_{\text{\tiny{F}}}=0.5$ with $\gamma=-1$. These two cases share the same Hubble parameter but have different dynamics. Note that with these choices, $\bar{\beta}_{\text{\tiny{F}}}$ is positive and zeros of $H_{,\phi}\propto\phi_{\text{\tiny{F}}}Ai(\phi_{\text{\tiny{F}}})$ are located at $\phi_{\text{\tiny{F}}}=0$ and isolated values of $\phi_{\text{\tiny{F}}}$ on negative real axis. Following the notation of reference [79], we show $s^{\text{th}}$ zeros of $Ai(z)$ and $Ai^{\prime}(z)$ with $a_{s}$ and $a^{\prime}_{s}$, respectively. Specifically, we will focus on the model around the $\phi_{\text{\tiny{F}}}=a_{1}\approx-2.34$ point where $H_{,\phi}=0$ and quasi-canonical behavior is expected. To avoid any ambiguity, we further restrict our discussion to the $a^{\prime}_{1}<\phi_{\text{\tiny{F}}}<a^{\prime}_{2}$ range, where the Hubble parameter remains positive for $M>0$. As we will demonstrate later, the inflaton naturally resides within this region, as inflation ends before the sign of Hubble parameter changes.
Refer to caption    Refer to captionRefer to caption     Refer to caption
@@
Figure 4: The plots of the Hubble parameter (4.31) (top left panel), number of $e$-folds (4.32) (top right panel), potential (4.33) and effective potential (4.35) (bottom left panel), and slow-roll parameters (4.36) and (4.37) (bottom right panel), once for $\beta_{\text{\tiny{F}}}=-0.5$ and $\gamma=1$ (solid lines) and once for $\beta_{\text{\tiny{F}}}=0.5$ and $\gamma=-1$ (dashed lines). The parameter $\mathcalboondox{V}_{0}$ in effective potential is taken to be $1.475$. Vertical grid lines in all panels correspond to the extrema of $H_{,\phi}$ at $\phi_{\text{\tiny{F}}}=a_{1}\approx-2.34$.
@@
The plot of $N(\phi_{\text{\tiny{F}}})$ can be seen in the top right panel of Figure 4. We see that $N(\phi_{\text{\tiny{F}}})$ is singular at $\phi_{\text{\tiny{F}}}=a_{1}$, where $H_{,\phi}\rightarrow 0$, as expected.
@@
where $\mathcalboondox{V}_{0}$ is the constant of integration. Potential (4.33) and effective potential (4.35) are plotted in Figure 4 (bottom left panel). The effective potential shows a roller-coaster profile with plateau-like regions connected to steep cliffs. Effective potential extrema are located at $a_{s}$ and $a^{\prime}_{s}$ which correspond to zeros of $H_{,\phi}$ and $H$, respectively. We recall that in FCR models the effective potential is not directly related to $\mathcal{F}$ and its concavity at extrema depends only on the value of the constant-roll parameter $\beta_{\text{\tiny{F}}}$. However, this is not the case for the potential. In Figure 4 we see that for $\beta_{\text{\tiny{F}}}=-0.5$ the effective potential is convex and for $\beta_{\text{\tiny{F}}}=0.5$ it is concave, while the potential is concave in both cases.
Depending on the field’s initial value, [^4] The dynamics of the constant propagation models can be completely determined by the field’s initial value as the field velocity is evaluated by equation (4.9) and we do not have freedom of choosing it. one can think of different inflationary scenarios. In the case of $\beta_{\text{\tiny{F}}}=-0.5$, the singularity of $N(\phi_{\text{\tiny{F}}})$ plot at $a_{1}$ shows that if the analytic field (that is described by constant $\beta_{\text{\tiny{F}}}$) starts at the neighbourhood of $\phi_{\text{\tiny{F}}}=a_{1}$, it will eventually rest at extrema. The field in $\beta_{\text{\tiny{F}}}=0.5$ case always rolls in the opposite direction to the $\beta_{\text{\tiny{F}}}=-0.5$ case.
Refer to caption     Refer to caption
@@
Figure 5: Another plot of the number of $e$-folds (4.32) (left panel), potential (4.33) (solid lines in right panel), and effective potential (4.35) (dashed lines in right panel) for $\gamma=1$ and $\mathcalboondox{V}_{0}=1.475$. Vertical grid line at $\phi_{\text{\tiny{F}}}=a_{1}$ corresponds to the singularity of function $N(\phi_{\text{\tiny{F}}})$.
Next, we derive and analyse slow-roll parameters. Using equations (4.9) and (4.31), for the first and second slow-roll parameters we find
@@
Figure 4 (bottom right panel) illustrates these parameters. At $\phi_{\text{\tiny{F}}}=a_{1}$, $\epsilon_{1}$ is approaching zero, while $\epsilon_{2}$ is roughly $2\beta_{\text{\tiny{F}}}$. When $\beta_{\text{\tiny{F}}}=-0.5$, the field moves toward the extrema of effective potential, where we observe quasi-canonical behavior, whereas when $\beta_{\text{\tiny{F}}}=0.5$, the field moves away from this point, which means we gradually leave the quasi-canonical phase. When $\beta_{\text{\tiny{F}}}$ is positive, $\epsilon_{1}$ grows in time and we can find points where the first slow-roll parameter reaches 1 and inflation ends. Additionally, since we have $\epsilon_{2}\approx 2\beta_{\text{\tiny{F}}}$ when $\phi_{\text{\tiny{F}}}\rightarrow a_{1}$, choosing $\beta_{\text{\tiny{F}}}<-1$ can potentially lead to a burst in the power spectrum and the formation of primordial black holes. We will talk more about this mechanism in another example in section 5.2.
@@
In Figure 4 we showed how the concavity of potential and effective potential changes around $\beta_{\text{\tiny{F}}}=0$. In Figure 5 (right panel) we plotted the potential and effective potential again, this time for $\beta_{\text{\tiny{F}}}=-2$, $-3$, and $-4$, around the critical point $\phi_{\text{\tiny{F}}}=a_{1}$, to show how the concavity changes around $\beta_{\text{\tiny{F}}}=-3$. We can see that the potential is convex for $\beta_{\text{\tiny{F}}}=-4$ and concave for $\beta_{\text{\tiny{F}}}=-2$. In each case, since $\mathcal{F}$ is negative, the concavity of effective potential is in the opposite direction of the potential. When $\beta_{\text{\tiny{F}}}=-3$, we have $\mathcalboondox{V}_{,\phi\phi}=\mathcalboondox{V}_{,\phi}=0$ and effective potential is flat throughout. There is, however, only one point (at the singularity of $N(\phi_{\text{\tiny{F}}})$) where the first and second derivatives of potential are zero. Actually, we do not have a flat USR-like potential for any choice of the constant-roll parameter $\beta_{\text{\tiny{F}}}$. To study the USR limit of this model, one should use the PCR condition or equivalently the HCR condition.